1. Inflation erodes the purchasing power of a nation's currency. T/F

2. Use the following table to answer the next two questions.

In Terms of USD

Country

Bid

Ask

Brazil

$.3033

$.3055

Euro

$1.2216

$1.2220

2-a. What is the ask quote for Brazilian reais (R$) in terms of euros (€)? a. €.2482 b. €3.9984 c. €.2501 d. €4.0290

2-b. What is the ask quote for euros in terms of reis? Round intermediate steps to four decimals.

a. R$4.0290

b. R$3.9987

c. R$.2482

d. R$.2501

e. None of the above

3. Indirect foreign intervention tends to be less effective than direct foreign intervention. T/F

4. Suppose that Australia wishes to use sterilized market intervention to increase the value of the Australian dollar. Australia's central bank would ____ Australian dollars and simultaneously_____Australian treasury securities.

a. buy, buy

b. sell, sell

c. buy, sell

d. sell, buy

e. Cannot be determined

5. One disadvantage of the eurozone is that member nations cannot create their own, independent monetary policy strategy to adjust for macroeconomic changes. T/F
Answer #1

1. Inflation erodes the purchasing power of a nation's currency. T/F - False

2. This is case of cross exchange rate

( R/US$ )Bid / (E/US$)ASK = E0.2482 option a

3 . ask quote for euros in terms of reis = 1/ 0.2482 = R$4.0290 option a

4. option c or Option d ....

5. True